38camcoo:

The average score at Laila's favorite golf course is 91, with a standard deviation of 10. In their most recent round at the golf course, Laila scored an 85. Calculate the z-score for Laila's round of golf.

luhivqqcherry:

Based on the given information, we can calculate the z-score for Laila's round of golf using the formula: z = (x - μ) / σ where: x = Laila's score = 85 μ = mean score = 91 σ = standard deviation = 10 Substituting the values, we get: z = (85 - 91) / 10 z = -0.6 Therefore, the z-score for Laila's round of golf is -0.6.

Jumper:

Step 1. Grab a calculator Step 2. Solve Use this z = (x - μ) / σ, where x is the individual score, μ is the mean score, and σ is the standard deviation. 😉 Laila's score (x) is 85, the mean score (μ) is 91, and the standard deviation (σ) is 10. Plug this into your calculator. If you don't get this answer, you're wrong. 👇 z = (85 - 91) / 10 Simplify it: z = -6 / 10 Answer is -0.6. 👑 (No GPT version).